40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Message ID: 2118     Entry time: Mon Mar 5 22:55:37 2018
Author: Craig, awade 
Type: DailyProgress 
Category: BEAT 
Subject: North and South Cavities Relocked 

While Craig was messing around on computers all day, awade got to work on the optics table aligning the North path.  He managed to lock the North at 60% visibility without even touching our new mode matching lens positions.  We think we can do better in the near future.

But while we had two TEM00 modes, we decided to get a betnote measurement.
Beatnote strength: +1 dBm (This is including an ND filter on our optics)
Beatnote frequency: 103.3 MHz

Also, apparently our new vaccan temp control PID script had a sign flip in it, so we've been heating our can maximally for a while today, up to 40 degrees C.  We fixed this, which will cause our can to go to 30 C and the beatnote will slew violently.  This made getting a beatnote ASD difficult. 

Attachment 1: NorthCavityAndSouthCavityTogetherInPerfectHarmony.jpg  2.229 MB  | Hide | Hide all
NorthCavityAndSouthCavityTogetherInPerfectHarmony.jpg
Attachment 2: SlowVoltsAndPIDSettings.jpg  3.202 MB  | Hide | Hide all
SlowVoltsAndPIDSettings.jpg
Attachment 3: StitchedSpectrum_TransBeatnote_FMDevn_10kHz_SR560Gain_20_Avgs_20_Span_102p4kHz_05-03-2018_231449_Spectrum.pdf  174 kB  | Hide | Hide all
StitchedSpectrum_TransBeatnote_FMDevn_10kHz_SR560Gain_20_Avgs_20_Span_102p4kHz_05-03-2018_231449_Spectrum.pdf
Attachment 4: StitchedSpectrum_TransBeatnote_FMDevn_10kHz_SR560Gain_20_Avgs_20_Span_102p4kHz_05-03-2018_231449.tgz  39 kB
ELOG V3.1.3-